OpenStudy (anonymous):

Perfect Square Equations 16x^2=9 I am not asking you to do the problem I just need help :)

OpenStudy (anonymous):

ok I will help

OpenStudy (anonymous):

Whenever you only have a square minus a number like this, the first thing to do is to check and see if you can take the square root of the number you are subtracting. If you can, then it makes things a bit easier. When you have to factor something where one square is subtracted from another square (9 = 3^2), it makes things easy. Here's the trick: you have to find the square root of each term. x^2 - y^2 = (x + y) (x - y) If you apply that basic formula to your problem, you need to find the square root of 16x^2 and of 9. Now, the square root of 16x^2 = 4x (because 4x*4x = 16x^2), and the square root of 9 = 3. Following the same guidelines above, if you put these two square roots together, where you add them in one set of parentheses and subtract them in the other set of parentheses, you will get the correct answer. (4x + 3) (4x - 3) = 16x^2 - 9 Since the two sets of parentheses are the same except for the plus and minus, the middle term that you would usually have cancels out. (4x + 3) (4x - 3) = 16x^2 + 12x - 12x - 9 (the 12x - 12x will cancel each other out and leave you with just...) 16x^2 - 9 (which is what you needed to factor)

OpenStudy (anonymous):



OpenStudy (anonymous):

Thank you @frankyp and @katragaddasaichandra :)

OpenStudy (anonymous):

ok give me the problem you want me to help out with

OpenStudy (anonymous):

5t^2 - 3= 7

OpenStudy (anonymous):

ok that equals t^2=2

OpenStudy (anonymous):

please vote for my design http://www.signazon.com/contest/grad2014/lkkajqe5/

OpenStudy (anonymous):

I got 35/5 @katragaddasaichandra ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

please vote for my design http://www.signazon.com/contest/grad2014/lkkajqe5/

OpenStudy (anonymous):

can u vote

OpenStudy (anonymous):

give me a badge

OpenStudy (anonymous):

now vote

OpenStudy (anonymous):

O.o erm. I already gave one to @katragaddasaichandra :(